# TUTORIAL ARTICLE Contraction Theory for Nonlinear Stability Analysis and Learning-based Control: A Tutorial Overview

### v7

@@ 1.2 Related Work
Recent applications of machine learning often consider challenging scenarios in the field of systems and control theory, where we only have access to system trajectory data generated by unknown underlying dynamics, and the assumptions in the aforementioned adaptive control techniques are no longer valid. For situations where the data is used for system identification of full/residual dynamics Shi et al. [2019], Singh et al. [2020], Manchester et al. [2021], Blocher et al. [2017] by a spectrally-normalized DNN Miyato et al. [2018], we can show by contraction theory that the model-based approaches (e.g. CV-STEM and NCM) are still utilizable to guarantee robustness against dynamics modeling errors and external disturbances. It is also proposed in Boffi et al. [2020] that we could directly learn certificate functions such as contraction metrics and their associated Lyapunov functions using trajectory data. Note that some of the theoretical results on gradient descent algorithms, essential in the field of data-driven machine learning, can be replaced by more general ones based on contraction and geodesical convexity Wensing and Slotine [2020].
1.3 Notation
For a square matrix $A^{n\times n}$, we use the notation $A\succ 0$, $A\succeq 0$, $A\prec 0$, and $A\preceq 0$ for the positive definite, positive semi-definite, negative definite, negative semi-definite matrices, respectively. The $\mathcal{L}_{p}$ norm in the extended space $\mathcal{L}_{pe}$ [Khalil, 2002, pp. 196-197], $p\in[1,\infty]$, is defined as $\|(y)_{\tau}\|_{\mathcal{L}_{p}}=\left(\int_{0}^{\tau}\|y(t)\|^{p}\right)^{{1}/{p}}<\infty$ for $p\in[1,\infty)$ and $\|(y)_{\tau}\|_{\mathcal{L}_{\infty}}=\sup_{t\geq 0}\|(y(t))_{\tau}\|<\infty$ for $p=\infty$, where $(y(t))_{\tau}$ is a truncation of $y(t)$, i.e., $(y(t))_{\tau}=0$ for $t>\tau$ and $(y(t))_{\tau}=y(t)$ for $0\leq t\leq\tau$ with $\tau\in\mathbb{R}_{\geq 0}$. Furthermore, we use $f_{x}=\partial f/\partial x$, $M_{x_{i}}=\partial M/\partial x_{i}$, and $M_{x_{i}x_{j}}=\partial^{2}M/(\partial x_{i}\partial x_{j})$, where $x_{i}$ and $x_{j}$ ate the $i$th and $j$th elements of $x\in\mathbb{R}^{n}$, for describing partial derivatives in a limited space. The other notations are given in Table 2.
@@ 1.3 Notation
2 Contraction Theory
We present a brief review of the results from Lohmiller and Slotine [1998], Slotine [2003], Wang and Slotine [2005], Pham et al. [2009], Chung et al. [2013], Tsukamoto and Chung [2021], Dani et al. [2015], Tsukamoto and Chung [2019]. They will be extensively used to provide formal robustness and stability guarantees for a variety of systems in the subsequent sections, simplifying and generalizing Lyapunov theory.
2.1 Fundamentals
@@
where $t\in\mathbb{R}_{\geq 0}$ is time, ${x}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n}$ the system state, and ${f}:\mathbb{R}^{n}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n}$ is a smooth function. Note that the smoothness of $f(x,t)$ guarantees existence and uniqueness of the solution to (2.1) for a given $x(0)=x_{0}$ at least locally [Khalil, 2002, pp. 88-95].
@@
A differential displacement, $\delta{x}$, is defined as an infinitesimal displacement at a fixed time as used in the calculus of variation [Kirk, 2004, p. 107], and (2.1) yields the following differential dynamics:
@@
Let us first present a special case of the comparison lemma [Khalil, 2002, pp. 102-103, pp. 350-353] to be used extensively throughout this paper.
@@
See [Khalil, 2002, pp. 659-660]. $\qed$
2.1.1 Contraction Theory and Contraction Metric
In Lyapunov theory, nonlinear stability of (2.1) is studied by constructing a Lyapunov function $V(x,t)$, one example of which is $V=x^{\top}P(x,t)x$. However, finding $V(x,t)$ for general nonlinear systems is challenging as $V(x,t)$ can be any scalar function of $x$ (e.g., a candidate $V(x,t)$ can be obtained by solving a PDE [Khalil, 2002, p. 211]). In contrast, as summarized in Table 1, contraction theory uses a differential Lyapunov function that is always a quadratic function of $\delta x$, i.e., $V(x,\delta x,t)=\delta x^{\top}M(x,t)\delta x$, thereby characterizing a necessary and sufficient condition for incremental exponential convergence of the multiple nonlinear system trajectories to one single trajectory. Thus, the problem of finding $V$ for stability analysis boils down to finding a finite-dimensional positive-definite matrix $M$, as illustrated in Figure 1 Lohmiller and Slotine [1998]. These properties to be derived in Theorem 2.1, which hold both for autonomous (i.e. time-invariant) and non-autonomous systems, epitomize significant methodological simplifications of stability analysis in contraction theory.
Figure 1: Lyapunov theory and contraction theory, where $V$ is a Lyapunov function and $\delta z=\Theta(x,t)\delta x$ for $M(x,t)=\Theta(x,t)\Theta(x,t)^{\top}\succ 0$ that defines a contraction metric (see Theorem 2.1).
@@ 2.1.1 Contraction Theory and Contraction Metric
If there exists a uniformly positive definite matrix ${M}({x},t)={{\Theta}}({x},t)^{\top}{{\Theta}}({x},t)\succ 0,~\forall x,t$, where $\Theta(x,t)$} defines a smooth coordinate transformation of $\delta x$, i.e., $\delta{z}={\Theta}(x,t)\delta{x}$, s.t. either of the following equivalent conditions holds for $\exists\alpha\in\mathbb{R}_{>0}$, $\forall x,t$:
@@
where the arguments $(x,t)$ of $M(x,t)$ and $\Theta(x,t)$ are omitted for notational simplicity, then all the solution trajectories of (2.1) converge to a single trajectory exponentially fast regardless of their initial conditions (i.e., contracting, see Definition 2.3), with an exponential convergence rate $\alpha$. The converse also holds.
@@
The proof of this theorem can be found in Lohmiller and Slotine [1998], but here we emphasize the use of the comparison lemma given in Lemma 2.1. Taking the time-derivative of a differential Lyapunov function of $\delta x$ (or $\delta z$), $V=\delta{z}^{\top}\delta{z}=\delta{x}^{\top}M(x,t)\delta{x}$, using the differential dynamics (2.2), we have
@@
where the conditions (2.5) and (2.6) are used with the generalized Jacobian $F$ in (2.5) obtained from $\delta\dot{z}=\dot{\Theta}\delta x+\Theta\delta\dot{x}=F\delta z$. We get $d\|\delta{z}\|/dt\leq-\alpha\|\delta{z}\|$ by $d\|\delta{z}\|^{2}/dt=2\|\delta{z}\|d\|\delta{z}\|/dt$, which then yields $\|\delta{z}(t)\|\leq\|\delta{z}_{0}\|e^{-\alpha t}$ by the comparison lemma of Lemma 2.1. Hence, any infinitesimal length $\|\delta{z}(t)\|$ and $\|\delta{x}(t)\|$, as well as $\delta z$ and $\delta x$, tend to zero exponentially fast. By path integration (see Definition 2.2 and Theorem 2.3), this immediately implies that the length of any finite path converges exponentially to zero from any initial conditions.
@@
which indeed implies that $\Xi\succeq\mathrm{I}\succ 0$ as (2.9) holds for any $\delta x$. Thus, $\Xi$ satisfies the contraction condition (2.6), i.e., $\Xi$ defines a contraction metric (see Definition 2.3Lohmiller and Slotine [1998]. $\qed$
@@
Since $M$ of Theorem 2.1 is positive definite, i.e., $v^{\top}Mv=\|\Theta v\|^{2}\geq 0,~\forall v\in\mathbb{R}^{n}$, and $\|\Theta v\|^{2}=0$ if and only if $v=0$, the equation $\Theta v=0$ only has a trivial solution $v=0$. This implies that $\Theta$ is always non-singular (i.e., $\Theta(x,t)^{-1}$ always exists).
@@
Let $\xi_{0}(t)$ and $\xi_{1}(t)$ denote some solution trajectories of (2.1). We say that (2.1) is incrementally exponentially stable if $\exists C,\alpha>0$ s.t. the following holds Angeli [2002]:
@@
for any $\xi_{0}(t)$ and $\xi_{1}(t)$. Note that, since we have $\|\xi_{1}(t)-\xi_{0}(t)\|=\|\int_{\xi_{0}}^{\xi_{1}}\delta x\|$ (see Theorem 2.3), Theorem 2.1 implies incremental stability of the system (2.1).
@@
The system (2.1) satisfying the conditions in Theorem 2.1 is said to be contracting, and a uniformly positive definite matrix $M$ that satisfies (2.6) defines a contraction metric. As to be discussed in Theorem 2.3 of Sec. 2.2, a contracting system is incrementally exponentially stable in the sense of Definition 2.2.
@@
One of the distinct features of contraction theory in Theorem 2.1 is incremental stability with exponential convergence. Consider the example given in Jouffroy and Slotine [2004]:
@@
and the contraction condition (2.6) for (2.12) can no longer be proven by $V=\|\delta x\|^{2}/2$, due to the lack of the skew-symmetric property of (2.11) in (2.12). This difficulty illustrates the difference between Lyapunov theory and contraction theory, where the former considers stability of (2.11) with respect to the equilibrium point, while the latter analyzes exponential convergence of any couple of trajectories in (2.11) with respect to each other (i.e., incremental stability in Definition 2.2Lohmiller and Slotine [1998], Jouffroy and Slotine [2004].
@@
Contraction defined by Theorem 2.1 guarantees incremental stability of their solution trajectories but does not require the existence of stable fixed points. Let us consider the following system for $x:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$:
@@
Using the constraint (2.6) of Theorem 2.1, we can easily verify that (2.13) is contracting as $M=\mathrm{I}$ defines its contraction metric with the contraction rate $\alpha=1$. However, since (2.13) has $x(t)=e^{t}/2+(x(0)-1/2)e^{-t}$ as its unique solution, it is not stable with respect to any fixed point.
@@
Consider a Linear Time-Invariant (LTI) system, $\dot{x}=f(x)=Ax$. Lyapunov theory states that the origin is globally exponentially stable if and only if there exists a constant positive-definite matrix $P\in\mathbb{R}^{n\times n}$ s.t. [Rugh, 1996, pp. 67-68]
@@
The contraction condition (2.6) can thus be viewed as a generalization of the Lyapunov stability condition (2.14) (see the generalized Krasovskii’s theorem [Slotine and Li, 1991, pp. 83-86]) in a nonlinear non-autonomous system setting, expressed in the differential formulation that permits a non-constant metric and pure differential coordinate change Lohmiller and Slotine [1998]. Furthermore, if $f(x,t)=A(t)x$ and $M=\mathrm{I}$, (2.6) results in $A(t)+A(t)^{\top}\preceq-2\alpha\mathrm{I}$ (i.e., all the eigenvalues of the symmetric matrix $A(t)+A(t)^{\top}$ remain strictly in the left-half complex plane), which is a known sufficient condition for stability of Linear Time-Varying (LTV) systems [Slotine and Li, 1991, pp. 114-115].
@@
Most of the learning-based techniques involving neural networks are based on optimizing their hyperparameters by gradient descent Robbins and Monro [1951]. Contraction theory provides a generalized view on the analysis of such continuous-time gradient-based optimization algorithms Wensing and Slotine [2020].
Let us consider a twice differentiable scalar output function $f:\mathbb{R}^{n}\times\mathbb{R}\rightarrow\mathbb{R}$, a matrix-valued function $M:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times n}$ with $M(x)\succ 0,~\forall x\in\mathbb{R}^{n}$, and the following natural gradient system Amari [1998]:
@@
Then, $f$ is geodesically $\alpha$-strongly convex for each $t$ in the metric defined by $M(x)$ (i.e., $H(x)\succeq\alpha M(x)$ with $H(x)$ being the Riemannian Hessian matrix of $f$ with respect to $M$ Udriste [1994]), if and only if (2.15) is contracting with rate $\alpha$ in the metric defined by $M$ as in (2.6) of Theorem 2.1, where $A={\partial h}/{\partial x}$. More specifically, the Riemannian Hessian verifies $H(x)=-(\dot{M}+MA+A^{\top}M)/2$. See Wensing and Slotine [2020] for details.
@@
Theorem 2.1 can be applied to other vector norms of $\|\delta{z}\|_{p}$ with, e.g., $p=1$ or $p=\infty$ Lohmiller and Slotine [1998]. It can also be shown that for a contracting autonomous system of the form $\dot{x}={f}({x})$, all trajectories converge to an equilibrium point exponentially fast.
2.1.2 Partial Contraction
Although satisfying the condition (2.6) of Theorem 2.1 guarantees exponential convergence of any couple of trajectories in (2.1), proving their incremental stability with respect to a subset of these trajectories possessing a specific property could be sufficient for some cases Tsukamoto and Chung [2021, 2021], Tsukamoto et al. [2021], Tsukamoto and Chung [2019], leading to the concept of partial contraction Wang and Slotine [2005].

### v8

@@ 1.2. Related Work
Recent applications of machine learning often consider challenging scenarios in the field of systems and control theory, where we only have access to system trajectory data generated by unknown underlying dynamics, and the assumptions in the aforementioned adaptive control techniques are no longer valid. For situations where the data is used for system identification of full/residual dynamics neurallander, 47710, cdc_systemid, 7992901 by a spectrally-normalized DNN miyato2018spectral, we can show by contraction theory that the model-based approaches (e.g. CV-STEM and NCM) are still utilizable to guarantee robustness against dynamics modeling errors and external disturbances. It is also proposed in boffi2020learning that we could directly learn certificate functions such as contraction metrics and their associated Lyapunov functions using trajectory data. Note that some of the theoretical results on gradient descent algorithms, essential in the field of data-driven machine learning, can be replaced by more general ones based on contraction and geodesical convexity beyondconvexity.
1.3. Notation
For a square matrix $A\in\mathbb{R}^{n\times n}$, we use the notation $A\succ 0$, $A\succeq 0$, $A\prec 0$, and $A\preceq 0$ for the positive definite, positive semi-definite, negative definite, and negative semi-definite matrices, respectively. The $\mathcal{L}_{p}$ norm in the extended space $\mathcal{L}_{pe}$ [Khalil:1173048, pp. 196-197], $p\in[1,\infty]$, is defined as $\|(y)_{\tau}\|_{\mathcal{L}_{p}}=\left(\int_{0}^{\tau}\|y(t)\|^{p}\right)^{{1}/{p}}<\infty$ for $p\in[1,\infty)$ and $\|(y)_{\tau}\|_{\mathcal{L}_{\infty}}=\sup_{t\geq 0}\|(y(t))_{\tau}\|<\infty$ for $p=\infty$, where $(y(t))_{\tau}$ is a truncation of $y(t)$, i.e., $(y(t))_{\tau}=0$ for $t>\tau$ and $(y(t))_{\tau}=y(t)$ for $0\leq t\leq\tau$ with $\tau\in\mathbb{R}_{\geq 0}$. Furthermore, we use $f_{x}=\partial f/\partial x$, $M_{x_{i}}=\partial M/\partial x_{i}$, and $M_{x_{i}x_{j}}=\partial^{2}M/(\partial x_{i}\partial x_{j})$, where $x_{i}$ and $x_{j}$ ate the $i$th and $j$th elements of $x\in\mathbb{R}^{n}$, for describing partial derivatives in a limited space. The other notations are given in Table 2.
@@ 1.3. Notation
2. Contraction Theory
We present a brief review of the results from Ref:contraction1, Ref:contraction2, Ref:contraction3, Pham2009, Ref:phasesync, mypaperTAC, Ref:Stochastic, mypaper. They will be extensively used to provide formal robustness and stability guarantees for a variety of systems in the subsequent sections, simplifying and generalizing Lyapunov theory.
2.1. Fundamentals
@@
where $t\in\mathbb{R}_{\geq 0}$ is time, ${x}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n}$ the system state, and ${f}:\mathbb{R}^{n}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n}$ is a smooth function. Note that the smoothness of $f(x,t)$ guarantees existence and uniqueness of the solution to (2.1) for a given $x(0)=x_{0}$ at least locally [Khalil:1173048, pp. 88-95].
@@
Differential Dynamics A differential displacement, $\delta{x}$, is defined as an infinitesimal displacement at a fixed time as used in the calculus of variation [citeulike:802300, p. 107], and (2.1) yields the following differential dynamics:
@@
Let us first present a special case of the comparison lemma [Khalil:1173048, pp. 102-103, pp. 350-353] to be used extensively throughout this paper.
@@
See [Khalil:1173048, pp. 659-660]. $\qed$
2.1.1. Contraction Theory and Contraction Metric
In Lyapunov theory, nonlinear stability of (2.1) is studied by constructing a Lyapunov function $V(x,t)$, one example of which is $V=x^{\top}P(x,t)x$. However, finding $V(x,t)$ for general nonlinear systems is challenging as $V(x,t)$ can be any scalar function of $x$ (e.g., a candidate $V(x,t)$ can be obtained by solving a PDE [Khalil:1173048, p. 211]). In contrast, as summarized in Table 1, contraction theory uses a differential Lyapunov function that is always a quadratic function of $\delta x$, i.e., $V(x,\delta x,t)=\delta x^{\top}M(x,t)\delta x$, thereby characterizing a necessary and sufficient condition for incremental exponential convergence of the multiple nonlinear system trajectories to one single trajectory. Thus, the problem of finding $V$ for stability analysis boils down to finding a finite-dimensional positive-definite matrix $M$, as illustrated in Figure 1 Ref:contraction1. These properties to be derived in Theorem LABEL:THM:Thm:contraction, which hold both for autonomous (i.e. time-invariant) and non-autonomous systems, epitomize significant methodological simplifications of stability analysis in contraction theory.
Figure 1: Lyapunov theory and contraction theory, where $V$ is a Lyapunov function and $\delta z=\Theta(x,t)\delta x$ for $M(x,t)=\Theta(x,t)\Theta(x,t)^{\top}\succ 0$ that defines a contraction metric (see Theorem LABEL:THM:Thm:contraction).
@@ 2.1.1. Contraction Theory and Contraction Metric
Deterministic ContractionThm:contraction If there exists a uniformly positive definite matrix ${M}({x},t)={{\Theta}}({x},t)^{\top}{{\Theta}}({x},t)\succ 0,~\forall x,t$, where $\Theta(x,t)$} defines a smooth coordinate transformation of $\delta x$, i.e., $\delta{z}={\Theta}(x,t)\delta{x}$, s.t. either of the following equivalent conditions holds for $\exists\alpha\in\mathbb{R}_{>0}$, $\forall x,t$:
@@
where the arguments $(x,t)$ of $M(x,t)$ and $\Theta(x,t)$ are omitted for notational simplicity, then all the solution trajectories of (2.1) converge to a single trajectory exponentially fast regardless of their initial conditions (i.e., contracting, see Definition LABEL:DEF:Def:contraction), with an exponential convergence rate $\alpha$. The converse also holds.
@@
The proof of this theorem can be found in Ref:contraction1, but here we emphasize the use of the comparison lemma given in Lemma 2.1. Taking the time-derivative of a differential Lyapunov function of $\delta x$ (or $\delta z$), $V=\delta{z}^{\top}\delta{z}=\delta{x}^{\top}M(x,t)\delta{x}$, using the differential dynamics (2.2), we have
@@
where the conditions (2.5) and (2.6) are used with the generalized Jacobian $F$ in (2.5) obtained from $\delta\dot{z}=\dot{\Theta}\delta x+\Theta\delta\dot{x}=F\delta z$. We get $d\|\delta{z}\|/dt\leq-\alpha\|\delta{z}\|$ by $d\|\delta{z}\|^{2}/dt=2\|\delta{z}\|d\|\delta{z}\|/dt$, which then yields $\|\delta{z}(t)\|\leq\|\delta{z}_{0}\|e^{-\alpha t}$ by the comparison lemma of Lemma 2.1. Hence, any infinitesimal length $\|\delta{z}(t)\|$ and $\|\delta{x}(t)\|$, as well as $\delta z$ and $\delta x$, tend to zero exponentially fast. By path integration (see Definition LABEL:DEF:Def:incremental and Theorem LABEL:THM:Thm:path_integral), this immediately implies that the length of any finite path converges exponentially to zero from any initial conditions.
@@
which indeed implies that $\Xi\succeq\mathrm{I}\succ 0$ as (2.9) holds for any $\delta x$. Thus, $\Xi$ satisfies the contraction condition (2.6), i.e., $\Xi$ defines a contraction metric (see Definition LABEL:DEF:Def:contractionRef:contraction1. $\qed$
@@
Since $M$ of Theorem LABEL:THM:Thm:contraction is positive definite, i.e., $v^{\top}Mv=\|\Theta v\|^{2}\geq 0,~\forall v\in\mathbb{R}^{n}$, and $\|\Theta v\|^{2}=0$ if and only if $v=0$, the equation $\Theta v=0$ only has a trivial solution $v=0$. This implies that $\Theta$ is always non-singular (i.e., $\Theta(x,t)^{-1}$ always exists).
@@
Incremental Exponential StabilityDef:incremental Let $\xi_{0}(t)$ and $\xi_{1}(t)$ denote some solution trajectories of (2.1). We say that (2.1) is incrementally exponentially stable if $\exists C,\alpha>0$ s.t. the following holds 989067:
@@
for any $\xi_{0}(t)$ and $\xi_{1}(t)$. Note that, since we have $\|\xi_{1}(t)-\xi_{0}(t)\|=\|\int_{\xi_{0}}^{\xi_{1}}\delta x\|$ (see Theorem LABEL:THM:Thm:path_integral), Theorem LABEL:THM:Thm:contraction implies incremental stability of the system (2.1).
@@
Contracting SystemsDef:contraction The system (2.1) satisfying the conditions in Theorem LABEL:THM:Thm:contraction is said to be contracting, and a uniformly positive definite matrix $M$ that satisfies (2.6) defines a contraction metric. As to be discussed in Theorem LABEL:THM:Thm:path_integral of Sec. 2.2, the metric induces the Riemannian distance, and a contracting system is incrementally exponentially stable in the sense of Definition LABEL:DEF:Def:incremental.
@@
Lyapunov Theory vs. Contraction Theory Iex:cont_ex1 One of the distinct features of contraction theory in Theorem LABEL:THM:Thm:contraction is incremental stability with exponential convergence. Consider the example given in Ref:contraction6:
@@
and the contraction condition (2.6) for (2.12) can no longer be proven by $V=\|\delta x\|^{2}/2$, due to the lack of the skew-symmetric property of (2.11) in (2.12). This difficulty illustrates the difference between Lyapunov theory and contraction theory, where the former considers stability of (2.11) with respect to the equilibrium point, while the latter analyzes exponential convergence of any couple of trajectories in (2.11) with respect to each other (i.e., incremental stability in Definition LABEL:DEF:Def:incrementalRef:contraction1, Ref:contraction6.
@@
Lyapunov Theory vs. Contraction Theory II Contraction defined by Theorem LABEL:THM:Thm:contraction guarantees incremental stability of their solution trajectories but does not require the existence of stable fixed points. Let us consider the following system for $x:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$:
@@
Using the condition (2.6) of Theorem LABEL:THM:Thm:contraction, we can easily verify that (2.13) is contracting as $M=\mathrm{I}$ defines its contraction metric with the contraction rate $\alpha=1$. However, since (2.13) has $x(t)=e^{t}/2+(x(0)-1/2)e^{-t}$ as its unique solution, it is not stable with respect to any fixed point.
@@
Contraction in Linear Time-Invariant Systemsex:lyapunov_contraction Consider a Linear Time-Invariant (LTI) system, $\dot{x}=f(x)=Ax$. Lyapunov theory states that the origin is globally exponentially stable if and only if there exists a constant positive-definite matrix $P\in\mathbb{R}^{n\times n}$ s.t. [lssbook, pp. 67-68]
@@
The contraction condition (2.6) can thus be viewed as a generalization of the Lyapunov stability condition (2.14) (see the generalized Krasovskii’s theorem [Ref_Slotine, pp. 83-86]) in a nonlinear non-autonomous system setting, expressed in the differential formulation that permits a non-constant metric and pure differential coordinate change Ref:contraction1. Furthermore, if $f(x,t)=A(t)x$ and $M=\mathrm{I}$, (2.6) results in $A(t)+A(t)^{\top}\preceq-2\alpha\mathrm{I}$ (i.e., all the eigenvalues of the symmetric matrix $A(t)+A(t)^{\top}$ remain strictly in the left-half complex plane), which is a known sufficient condition for stability of Linear Time-Varying (LTV) systems [Ref_Slotine, pp. 114-115].
@@
Contraction in Gradient Descent Most of the learning-based techniques involving neural networks are based on optimizing their hyperparameters by gradient descent sgd. Contraction theory provides a generalized view on the analysis of such continuous-time gradient-based optimization algorithms beyondconvexity.
Let us consider a twice differentiable scalar output function $f:\mathbb{R}^{n}\times\mathbb{R}\rightarrow\mathbb{R}$, a matrix-valued function $M:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times n}$ with $M(x)\succ 0,~\forall x\in\mathbb{R}^{n}$, and the following natural gradient system doi:10.1162/089976698300017746:
@@
Then, $f$ is geodesically $\alpha$-strongly convex for each $t$ in the metric defined by $M(x)$ (i.e., $H(x)\succeq\alpha M(x)$ with $H(x)$ being the Riemannian Hessian matrix of $f$ with respect to $M$ gconvex), if and only if (2.15) is contracting with rate $\alpha$ in the metric defined by $M$ as in (2.6) of Theorem LABEL:THM:Thm:contraction, where $A={\partial h}/{\partial x}$. More specifically, the Riemannian Hessian verifies $H(x)=-(\dot{M}+MA+A^{\top}M)/2$. See beyondconvexity for details.
@@
Theorem LABEL:THM:Thm:contraction can be applied to other vector norms of $\|\delta{z}\|_{p}$ with, e.g., $p=1$ or $p=\infty$ Ref:contraction1. It can also be shown that for a contracting autonomous system of the form $\dot{x}={f}({x})$, all trajectories converge to an equilibrium point exponentially fast.
2.1.2. Partial Contraction
Although satisfying the condition (2.6) of Theorem LABEL:THM:Thm:contraction guarantees exponential convergence of any couple of trajectories in (2.1), proving their incremental stability with respect to a subset of these trajectories possessing a specific property could be sufficient for some cases mypaperTAC, ncm, nscm, mypaper, leading to the concept of partial contraction Ref:contraction3.